please help me solve: 5/8x+1/16x=9/16+x
the solution is x=?
*these are fractions is the problem*

Answer by rfer(16322)

5/8x+1/16x=9/16+x
10/16x+1/16x=9/16+16/16x
11/16x-16/16x=9/16
-5/16x=9/16
divide by -5/16
x=(9/16)/(-5/16)
invert and multiply
x=9/16*-16/5
x=-144/80
x=-1 4/5